# Graph Reinforcement Learning for Combinatorial Optimization: A Survey and Unifying Perspective

### v1

@@ 3.3 Causal Discovery
Building a graph so as to maximize an objective function also has applications in causal inference and reasoning. In this area, Directed Acyclic Graphs (DAGs) are commonly used to represent statistical and causal dependencies between random variables. The problem of causal discovery asks to find the DAG that best explains the observed data, which can be quantified by score functions such as the Bayesian Information Criterion (BIC) (Schwarz, 1978).
@@
Zhu et al. (2020) proposed an encoder-decoder method that captures a probability of existence for each edge and is used to generate graphs in a one-shot fashion. RL-BIC is an actor-critic algorithm that also leverages the continuous characterization of acyclicity proposed by Zheng et al. (2018) in the reward function together with the score function itself to ensure that generated graphs do not contain cycles. The evaluation, conducted on synthetic random graphs and a benchmark in the biological domain, highlights the flexibility of the framework to accommodate varying data generation processes and score functions.
@@
Building on this work, Wang et al. (2021) considered carrying out the search in the space of orderings. An ordering is a permutation over the nodes that places a constraint on the possible relationships, i.e., for every node, its parents must come before it in the permutation. From the ordering, a causal graph can be constructed by (1) translating the ordering to a fully-connected DAG; and (2) pruning away statistically insignificant relationships. The proposed CORL method formulates actions as the selection of a node to add to the ordering and makes use of two reward signals, of which one is provided at the end of the episode when the graph is scored, and one is given after each decision-making step by exploiting the decomposability of BIC. The authors use an actor-critic RL algorithm with an encoder-decoder architecture and achieve better scalability than RL-BIC.
@@
Yang et al. (2023a) also leveraged the smaller search space of causal orderings and formulated actions as the selection of a node to add to the ordering. Instead of optimizing the score of a single source graph, the proposed RCL-OG method learns the posterior distribution of orderings given the observed data, basing this on the fact that the true DAG may not be identifiable in some settings. The resulting probabilistic model can be used for sampling orderings. An encoder-decoder architecture is used in conjunction with several Q-networks (one per layer of the order graph) trained using the DQN algorithm. Authors show that the correct posterior is accurately recovered in some simple examples and that better or comparable performance is obtained on a series of synthetic and real-world benchmarks.
@@
Darvariu et al. (2023b) is another work that approached the causal discovery problem with RL to search directly in the space of DAGs. Unlike RL-BIC, which performs one-shot graph generation, the authors formulated an MDP in which the graph is constructed edge-by-edge. The proposed CD-UCT method features an incremental algorithm for determining the candidate edges whose addition would cause the graph to become cyclic, which are excluded from the action space. The evaluation was carried out on benchmarks from the biological domain as well as synthetic graphs. CD-UCT was shown to substantially outperform RL-BIC even with orders of magnitude fewer score function evaluations, and also to compare favorably to Greedy Search.
@@ 3.4 Molecular Optimization
The final class of works we discuss in this section target molecular optimization. This is a fundamental task in chemistry, with applications in drug screening and material discovery. Molecules with various desirable properties such as drug-likeness and ease of synthesizability are sought. To represent molecules as graphs, atoms are mapped to nodes and edges to bonds.
You et al. (2018a) considered learning to construct molecular graphs. The objective functions that the method seeks to optimize are the drug-likeness and synthetic accessibility of molecules. The action space is defined as the addition of bonds or certain chemical substructures, and the transition function enables the environment to enforce validity rules with respect to physical laws. The proposed approach, called Graph Convolutional Policy Network (GCPN), uses GCN to compute the embeddings for state representation and trains the policy using PPO. In addition to the objective functions, the reward structure incentivizes the method to generate molecules that are similar to a given dataset of examples. GCPN was shown to outperform a series of previous generative models for the task.
@@ 4 Graph Process Optimization
Table 2: High-level summary of works that use Graph Reinforcement Learning for process optimization.
Problem
Decision Space
Objective Function
Base RL Algorithm Function Approximation Citation
Routing on Networks
Split ratios, edge weights used with softmin routing
Maximum Link Utilization
TRPO MLP (Valadarsky et al., 2017)
Split ratios
Delay, Throughput (multi-objective)
DDPG MLP (Xu et al., 2018b)
Flows to be rerouted with LP
Maximum Link Utilization
REINFORCE CNN (Zhang et al., 2020)
Selecting node as flow middlepoint
Maximum Link Utilization
PPO MPNN (Almasan et al., 2021)
Edge weights used with softmin routing
Maximum Link Utilization
PPO Graph Network (Hope & Yoneki, 2021)
Network Games
Continuous action vector used for edge formation
Recovering per-agent objective function
MA-AIRL MPNN (Trivedi & Zha, 2020)
Node to include in maximal independent set
Social Welfare, Fairness
UCT, GIL (Imitation Learning) S2V (Darvariu et al., 2021b)
Spreading Processes
Node to be influenced
Maximize influenced vertices
DQN GCN (Manchanda et al., 2020)
Node to be influenced (uncertain)
Maximize influenced vertices
DQN S2V (Chen et al., 2021)
Ranking over nodes
Minimize infected nodes, maximize influenced vertices
PPO 2 Custom GNNs (Meirom et al., 2021)
Search and Navigation
Next node towards unknown target
Identifying target node
REINFORCE LSTM (Das et al., 2018)
Next node towards unknown target
Identifying target node
UCT with Policy, Value Networks RNN (Shen et al., 2018)
Next cluster, next node towards unknown target
Identifying target node
REINFORCE LSTM (Zhang et al., 2022)
Next node towards known target
Number of node expansions until reaching target
SaIL (Imitation Learning) MLP (Bhardwaj et al., 2017)
Next node towards known target
Number of node expansions until reaching target
PHIL (Imitation Learning) Custom MPNN (Pándy et al., 2022)
Next node (no explicit target)
Minimize topological gaps, maximize compressibility
DQN GraphSAGE (Patankar et al., 2023)
@@ 4.1 Routing on Networks
Techniques for routing traffic across networks find applicability in a variety of scenarios including the Internet, road networks, and supply chains. [^2] We note that some works refer to problems in the TSP and VRP family as “routing” problems, since they involve devising routes, i.e., sequences of nodes to be visited. This is an unfortunate clash in terminology with problems that involve the routing of flows over graphs, which are discussed in this section. The problems have little in common in structure and solution methods beyond this superficial naming similarity. In the RL literature, routing across a network topology has been approached from two different perspectives. A first wave of interest considered routing at the packet level in a multi-agent RL formulation. The second, more recent, wave of interest originated in the computer networks community, which has begun to recognize the potential of ML methods in this space (Feamster & Rexford, 2017; Jiang et al., 2017). This work generally considers routing at the flow level rather than the more granular packet level, which tends to be a more scalable formulation of the problem, and is more aligned to current routing infrastructure.
@@ 4.2 Network Games
Network games (Jackson & Zenou, 2015) represent decentralized game-theoretic scenarios in which a social network connects self-interested agents. The actors have the liberty to take individual actions and derive utility based on their neighbors’ actions as well as their own. Typical questions of interest are finding equilibrium configurations of the game (from which agents would not unilaterally deviate given self-interested behavior) and devising strategies for incentivizing agents to move towards desirable configurations (e.g., as quantified by an objective function).
@@
Darvariu et al. (2021b) considered the problem of finding an equilibrium in the networked best-shot public goods game that optimizes an objective function. The authors exploited the correspondence of equilibria in this game with maximal independent sets – a set of nodes defined such that no two nodes are adjacent and each node is either a member of the set or not. The incremental construction of a maximal independent set was formulated as an MDP in which actions consist of node selections to be added to the set. Social welfare and fairness were considered as objectives. The authors leveraged the collection of demonstrations of an expert policy using Monte Carlo Tree Search, which are used by the proposed Graph Imitation Learning (GIL) technique for training a S2V-parameterized policy that generalizes to different problem instances and sizes. The method outperforms prior heuristics based on simulated annealing and a distributed search based on side payments.
@@ 4.3 Spreading Processes
Mathematical models of spreading processes are applicable for capturing the dynamics of phenomena such as the spread of disease, knowledge and innovation, or influence in a social network. They have a rich history in the mathematical study of epidemics, with models such as Susceptible, Infected, Recovered (SIR) enabling tractable analytical study (Kermack & McKendrick, 1927). Typical questions that are considered include the existence of a tipping point in the spread (beyond which the phenomenon propagates to the entire network) or how to best isolate nodes in the network to achieve containment. Several recent papers have considered applications of RL to spreading processes on graphs, with the techniques being applicable to more complex interaction mechanisms.
Manchanda et al. (2020) set out to improve the scalability of prior RL methods for combinatorial optimization problems such as S2V-DQN. The authors proposed to (1) train a model by Supervised Learning that predicts whether a node is likely to be part of the solution and (2) train a policy by RL that only operates on this subset of nodes. The GCN architecture was adopted and the model is trained with Q-learning. GCOMB was applied on the influence maximization problem, obtaining similar solution quality to prior methods while scaling to substantially larger graphs with billions of nodes.
@@ 4.4 Search and Navigation
Let us consider the first line of work, which addresses graph search in the context of reasoning in knowledge graphs. The task is typically formulated as completing a query: given an entity (e.g., Paul Erdős) and a relation (e.g., country of birth), the goal is to find the missing entity (e.g., Hungary). This is realized through guided walks over the knowledge graph. A model is trained using queries that are known to be true, and subsequently applied to tuples for which the knowledge is incomplete.
@@
A recent work by Zhang et al. (2022) addressed the issue of degrading performance of RL models for knowledge graph completion with increases in path length (e.g., MINERVA limits the path length to $3$ due to this). The authors proposed a hierarchical RL design with two policies that act cooperatively: one higher-level policy for picking the cluster in the knowledge graph to be searched, and a fine-grained policy that operates at the entity level. The initial clustering is performed using embeddings obtained with the TransE algorithm (Bordes et al., 2013) and K-means. The method outperforms MINERVA and M-Walk, particularly when answering queries over long paths.
@@
Bhardwaj et al. (2017) considered precisely the robotic motion planning use case, in which the graph corresponds to possible configurations of the robot, and edges are mapped to a set of valid maneuvers. The authors formulated this task as a POMDP and used a simple MLP to parameterize the policy, which was trained using imitation learning. The authors leveraged the existence of a powerful oracle algorithm whose computational cost prevents its use at runtime but may be queried during the training procedure. The optimization target is to minimize the expected difference between the Q-values (often referred to as cost-to-go in robotics, in which the equivalent goal is to minimize cost instead of maximizing reward) of the agent and the optimal Q-values supplied by the oracle. SaIL was shown to outperform several simple heuristics, an SL model, and model-free RL.

### v2

@@ 3.3 Causal Discovery
Building a graph so as to maximize an objective function also has applications in causal inference and reasoning. In this area, Directed Acyclic Graphs (DAGs) are commonly used to represent statistical and causal dependencies between random variables. Causal discovery can be approached as a combinatorial optimization problem that asks to find the DAG structure that best explains the observed data. This can be quantified by score functions such as the Bayesian Information Criterion (BIC) (Schwarz, 1978), which serve as objective functions in this context.
In this setting, Greedy Equivalence Search (Chickering, 2002) carries out a greedy search in the space of Markov equivalence classes. It is a cornerstone method and the most-cited score-based approach at the time of writing when compared to other such methods identified by a recent survey (Hasan et al., 2023). Another notable combinatorial technique for discrete random variables is GOBNILP (Cussens, 2011), an exact method that leverages an Integer Programming formulation of the problem. RL approaches have begun to be explored recently as a means of achieving flexibility in the data generation and score functions that can be used, as well as to improve on the simplistic greedy search mechanism.
@@
RL-BIC (Zhu et al., 2020) is an actor-critic algorithm that leverages the continuous characterization of acyclicity proposed by Zheng et al. (2018) in the reward function together with the score function itself to ensure that generated graphs do not contain cycles. The authors used an encoder-decoder model to approximate the policy. The encoder is based on the Transformer architecture, while the decoder is a single-layer non-linear transformation whose output is a square matrix used for predicting the probability of each possible edge in the graph. The evaluation, conducted on synthetic random graphs and a benchmark in the biological domain, highlights the flexibility of the framework to accommodate varying data generation processes and score functions.
@@
Building on this work, Wang et al. (2021) considered carrying out the search in the space of orderings. An ordering is a permutation over the nodes that places a constraint on the possible relationships, i.e., for every node, its parents must come before it in the permutation. From the ordering, a causal graph can be constructed by (1) translating the ordering to a fully-connected DAG; and (2) pruning away statistically insignificant relationships. The proposed CORL method formulates actions as the selection of a node to add to the ordering and makes use of two reward signals, of which one is provided at the end of the episode when the graph is scored, and one is given after each decision-making step by exploiting the decomposability of BIC. CORL uses function approximation for the policy through an encoder-decoder architecture. It features the same encoder as (Zhu et al., 2020) and an LSTM-based decoder that outputs an ordering node-by-node (with already-chosen nodes being masked). The authors use an actor-critic RL algorithm and achieve better scalability than RL-BIC.
@@
Yang et al. (2023a) also leveraged the smaller search space of causal orderings and formulated actions as the selection of a node to add to the ordering. Instead of optimizing the score of a single source graph, the proposed RCL-OG method learns the posterior distribution of orderings given the observed data, basing this on the fact that the true DAG may not be identifiable in some settings. The resulting probabilistic model can be used for sampling orderings. The method uses a symmetric learning architecture that leverages Transformers both for the encoder and the decoder. The output of the decoder is passed in input to an MLP to obtain estimates of the $Q$-values for each action (i.e., the next node to include in the ordering). The approach features several $Q$-networks (one per layer of the order graph) trained using the DQN algorithm. Authors show that the correct posterior is accurately recovered in some simple examples and that better or comparable performance is obtained on a series of synthetic and real-world benchmarks.
@@
Darvariu et al. (2023b) is another work that approached the causal discovery problem with RL to search directly in the space of DAGs. Unlike RL-BIC, which performs one-shot graph generation, the authors formulated an MDP in which the graph is constructed edge-by-edge. The proposed CD-UCT method features an incremental algorithm for determining the candidate edges whose addition would cause the graph to become cyclic, which are excluded from the action space. The evaluation was carried out on benchmarks from the biological domain as well as synthetic graphs. CD-UCT was shown to substantially outperform RL-BIC even with orders of magnitude fewer score function evaluations, and also to compare favorably to Greedy Search. The approach is based on MCTS and does not rely on function approximation.
@@ 3.4 Molecular Optimization
The final class of works we discuss in this section target molecular optimization. This is a fundamental task in chemistry, with applications in drug screening and material discovery. Molecules with various desirable properties such as drug-likeness and ease of synthesizability are sought. To represent molecules as graphs, atoms are mapped to nodes and edges to bonds. The catalogue of non-RL approaches to this problem is considerable. As discussed in the beginning of this section, sampling from a generative model trained on a molecule dataset has been considered. Other approaches include screening (i.e., performing an exhaustive search over) a database of known molecules; using metaheuristics such as hill climbing and genetic algorithms; and performing Bayesian optimization. We also note that RL methods operating on non-graph representations such as strings have also been proposed. For an extensive overview and experimental benchmark comparison of these methods, we refer the reader to the work of Gao et al. (2022). In the following, we review RL-based works that use a graph-based formulation of the molecular optimization problem.
You et al. (2018a) considered learning to construct molecular graphs using RL. The objective functions that the method seeks to optimize are the drug-likeness and synthetic accessibility of molecules. The action space is defined as the addition of bonds or certain chemical substructures, and the transition function enables the environment to enforce validity rules with respect to physical laws. The proposed approach, called Graph Convolutional Policy Network (GCPN), uses GCN to compute the embeddings for state representation and trains the policy using PPO. In addition to the objective functions, the reward structure incentivizes the method to generate molecules that are similar to a given dataset of examples. GCPN was shown to outperform a series of previous generative models for the task.
@@ 4 Graph Process Optimization
Table 1: High-level summary of works that use Graph Reinforcement Learning for structure optimization (Section 3).
Problem Decision Space Objective Function Base RL Algorithm Function Approximation Citation
Attacking GNNs Edge addition and removal Misclassification rate of GCN DQN S2V (Dai et al., 2018)
Edge addition for newly injected nodes Misclassification rate of GCN DQN S2V (Sun et al., 2020)
Degree-preserving edge rewiring Misclassification rate of GCN REINFORCE S2V (Ma et al., 2021)
Network Design Edge addition Resilience DQN S2V (Darvariu et al., 2021a)
Edge addition with spatial restrictions Resilience, Efficiency UCT (Darvariu et al., 2023a)
Edge rewiring Degree entropy, Maximum Entropy Random Walk DQN S2V (Doorman et al., 2022)
Degree-preserving edge rewiring Resilience, Efficiency PPO FireGNN (GIN-based) (Yang et al., 2023b)
Edge addition Recovering implicit objective MaxEnt Inverse RL MPNN (Trivedi et al., 2020)
Causal Discovery One-shot graph generation Bayesian Information Criterion Actor-Critic Encoder-Decoder (Zhu et al., 2020)
Node ordering Bayesian Information Criterion Actor-Critic Encoder-Decoder (Wang et al., 2021)
Node ordering Bayesian Information Criterion DQN Encoder-Decoder (Yang et al., 2023a)
Edge addition Bayesian Information Criterion UCT (Darvariu et al., 2023b)
Molecule Optimization Edge addition Drug-likeness, Synthetic accessibility PPO GCN (You et al., 2018a)
Node addition, Edge addition and removal Drug-likeness, Similarity to prior molecule DQN Molecular fingerprint (Zhou et al., 2019)
Table 2: High-level summary of works that use Graph Reinforcement Learning for process optimization (Section 4).
Problem Decision Space Objective Function Base RL Algorithm Function Approximation Citation
Routing on Networks Split ratios, edge weights used with softmin routing Maximum Link Utilization TRPO MLP (Valadarsky et al., 2017)
Split ratios Delay, Throughput (multi-objective) DDPG MLP (Xu et al., 2018b)
Flows to be rerouted with LP Maximum Link Utilization REINFORCE CNN (Zhang et al., 2020)
Selecting node as flow middlepoint Maximum Link Utilization PPO MPNN (Almasan et al., 2021)
Edge weights used with softmin routing Maximum Link Utilization PPO Graph Network (Hope & Yoneki, 2021)
Network Games Continuous action vector used for edge formation Recovering per-agent objective function MA-AIRL MPNN (Trivedi & Zha, 2020)
Node to include in maximal independent set Social Welfare, Fairness UCT, GIL (Imitation Learning) S2V (Darvariu et al., 2021b)
Spreading Processes Node to be influenced Maximize influenced vertices DQN GCN (Manchanda et al., 2020)
Node to be influenced (uncertain) Maximize influenced vertices DQN S2V (Chen et al., 2021)
Ranking over nodes Minimize infected nodes, Maximize influenced vertices PPO 2 Custom GNNs (Meirom et al., 2021)
Search and Navigation Next node towards unknown target Identifying target node REINFORCE LSTM (Das et al., 2018)
Next node towards unknown target Identifying target node UCT with Policy, Value Networks RNN (Shen et al., 2018)
Next cluster, next node towards unknown target Identifying target node REINFORCE LSTM (Zhang et al., 2022)
Next node towards known target Number of node expansions until reaching target SaIL (Imitation Learning) MLP (Bhardwaj et al., 2017)
Next node towards known target Number of node expansions until reaching target PHIL (Imitation Learning) Custom MPNN (Pándy et al., 2022)
Next node (no explicit target) Minimize topological gaps, Maximize compressibility DQN GraphSAGE (Patankar et al., 2023)
@@ 4.1 Routing on Networks
Techniques for routing traffic across networks find applicability in a variety of scenarios including the Internet, road networks, and supply chains. [^3] We note that some works refer to problems in the TSP and VRP family as “routing” problems, since they involve devising routes, i.e., sequences of nodes to be visited. This is an unfortunate clash in terminology with problems that involve the routing of flows over graphs, which are discussed in this section. The problems have little in common in structure and solution methods beyond this superficial naming similarity. In the RL literature, routing across a network topology has been approached from two different perspectives. A first wave of interest considered routing at the packet level in a multi-agent RL formulation. The second, more recent, wave of interest originated in the computer networks community, which has begun to recognize the potential of ML methods in this space (Feamster & Rexford, 2018; Jiang et al., 2017). This work generally considers routing at the flow level rather than the more granular packet level, which tends to be a more scalable formulation of the problem, and is more aligned to current routing infrastructure.
We now briefly discuss non-RL methods. The flow routing problem can be solved optimally by Linear Programming (Tardos, 1986). However, this requires the unrealistic assumption that the demands are known a priori and do not change (Fortz & Thorup, 2002). Shortest-path algorithms are also applicable, being commonly deployed in routing infrastructure via protocols, such as Open Shortest-Path First (OSPF) (Clark, 2003). The problem of determining weights for OSPF routing cannot be formulated as an LP and has been shown to be $\mathcal{NP}$-hard, motivating the use of local search heuristics (Fortz & Thorup, 2000). Simpler heuristics, such as setting weights inversely proportionally to capacity, have also been recommended by hardware manufacturers (Cisco, 2005).
@@ 4.2 Network Games
Network games (Jackson & Zenou, 2015) represent decentralized game-theoretic scenarios in which a social network connects self-interested agents. The actors have the liberty to take individual actions and derive utility based on their neighbors’ actions as well as their own. Typical questions of interest are finding equilibrium configurations of the game (from which agents would not unilaterally deviate given self-interested behavior) and devising strategies for incentivizing agents to move towards desirable configurations (e.g., as quantified by an objective function). In general, computing an equilibrium is often computationally challenging or even intractable (McKelvey & McLennan, 1996; Conitzer & Sandholm, 2003), motivating the use of approximate methods.
@@
Darvariu et al. (2021b) considered the problem of finding an equilibrium in the networked best-shot public goods game that optimizes an objective function. The authors exploited the correspondence of equilibria in this game with maximal independent sets – a set of nodes defined such that no two nodes are adjacent and each node is either a member of the set or not. The incremental construction of a maximal independent set was formulated as an MDP in which actions consist of node selections to be added to the set. Social welfare and fairness were considered as objectives. The authors leveraged the collection of demonstrations of an expert policy using MCTS, which are used by the proposed Graph Imitation Learning (GIL) technique for training a S2V-parameterized policy that generalizes to different problem instances and sizes. The method outperforms a number of prior approaches for this task including best-response dynamics, a method based on simulated annealing (Dall’Asta et al., 2011), a distributed search algorithm based on side payments (Levit et al., 2018), and heuristics that allocate resources based on local node properties.
@@ 4.3 Spreading Processes
Mathematical models of spreading processes are applicable for capturing the dynamics of phenomena such as the spread of disease, knowledge and innovation, or influence in a social network. They have a rich history in the mathematical study of epidemics, with models such as Susceptible, Infected, Recovered (SIR) enabling tractable analytical study (Kermack & McKendrick, 1927). Typical questions that are considered include the existence of a tipping point in the spread (beyond which the phenomenon propagates to the entire network) or how to best isolate nodes in the network to achieve containment.
The problem of influence maximization in a social network is $\mathcal{NP}$-hard and, for this reason, it has been approached with greedy search (Kempe et al., 2003) and hand-crafted heuristics (Liu et al., 2017), as exact methods are only usable at a small scale. For epidemic models, heuristics based on node properties (such as degree and betweenness) are typically used to identify nodes that would be influential in the spreading and hence should be isolated (Pastor-Satorras et al., 2015).
Several recent papers have considered applications of RL to spreading processes on graphs, with the techniques being applicable to more complex interaction mechanisms. Manchanda et al. (2020) set out to improve the scalability of prior RL methods for combinatorial optimization problems such as S2V-DQN. The authors proposed to (1) train a model by Supervised Learning that predicts whether a node is likely to be part of the solution and (2) train a policy by RL that only operates on this subset of nodes. The GCN architecture was adopted and the model is trained with Q-learning. GCOMB was applied on the influence maximization problem, obtaining similar solution quality to prior methods while scaling to substantially larger graphs with billions of nodes.
@@ 4.4 Search and Navigation
Let us consider the first line of work, which addresses graph search in the context of reasoning in knowledge graphs. The task is typically formulated as completing a query: given an entity (e.g., Paul Erdős) and a relation (e.g., country of birth), the goal is to find the missing entity (e.g., Hungary). This is realized through guided walks over the knowledge graph. A model is trained using queries that are known to be true, and subsequently applied to tuples for which the knowledge is incomplete. Notable traditional approaches for this task are TransE (Bordes et al., 2013) and TransR (Lin et al., 2015), which operate by embedding entities such that relations are interpreted as translations over the embedding space. The embeddings, which are learned by gradient descent, can subsequently be used to rank candidate entities for link prediction.
@@
A recent work by Zhang et al. (2022) addressed the issue of degrading performance of RL models for knowledge graph completion with increases in path length (e.g., MINERVA limits the path length to $3$ due to this). The authors proposed a hierarchical RL design with two policies that act cooperatively: one higher-level policy for picking the cluster in the knowledge graph to be searched, and a fine-grained policy that operates at the entity level. The initial clustering is performed using embeddings obtained with the TransE algorithm and K-means. The method outperforms MINERVA and M-Walk, particularly when answering queries over long paths.
@@
Bhardwaj et al. (2017) considered precisely the robotic motion planning use case, in which the graph corresponds to possible configurations of the robot, and edges are mapped to a set of valid maneuvers. The authors formulated this task as a POMDP and used a simple MLP to parameterize the policy, which was trained using imitation learning. The authors leveraged the existence of a powerful oracle algorithm whose computational cost prevents its use at runtime but may be queried during the training procedure. The optimization target is to minimize the expected difference between the Q-values (often referred to as cost-to-go in robotics, in which the equivalent goal is to minimize cost instead of maximizing reward) of the agent and the optimal Q-values supplied by the oracle. Search as Imitation Learning (SaIL) was shown to outperform several simple heuristics based on Euclidean and Manhattan distances, an SL model, and model-free RL.